OpenStudy (anonymous):

The graph of which equation has a vertex of (-2, 0)?

OpenStudy (anonymous):

after you post everything i will help you

OpenStudy (helder_edwin):

u forgot to post the equations!!

OpenStudy (anonymous):

A) y = (x - 2)2 + 2 B) y = 3(x + 2)2 + 1 C) y = 3(x - 2)2 D) y = 3(x + 2)2 @helder_edwin @FriedRice its been a long day so sorry

OpenStudy (anonymous):

do you know how to find the vertex from the vertex equation??

OpenStudy (anonymous):

this is the formula y = a(x-h)^2 + k the question is telling you to just plug in (h,k) or(-2,0)

OpenStudy (anonymous):

so y = a(x + 2)^2 + 0

OpenStudy (helder_edwin):

yes.

OpenStudy (anonymous):

do yo have the picture of the gragh because they have a slope and i dont know which one can be correct

OpenStudy (anonymous):

graph*

OpenStudy (helder_edwin):

which of the equations u gave us looks like what u just wrote?

OpenStudy (anonymous):

or replace \(x\) by \(-2\) (in your head) and see which one gives \(0\)

OpenStudy (anonymous):

i wasn't given a graph , just a list of equations @FriedRice

OpenStudy (helder_edwin):

u don't need the graph.

OpenStudy (anonymous):

D? @helder_edwin

OpenStudy (anonymous):

then the answer is D

OpenStudy (helder_edwin):

yes. very good.

OpenStudy (anonymous):

thank you all!